You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100409 (14 Apr 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 April08-09

Observer: A. Fontana, K.Boutsia, R. Speziali
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LUCIFER1

Summary

No clouds at the beginning of the night. Photometric night with good seeing (0.7-1.0)
Almost 2 hrs lost because of flipped MOS masks and unsuitable guiding stars.

Data Summary

Project Objects Filter ExpTime Notes
Grazian GOODS-N 200H+K OrderSep 2hrs arc - flat - tell
Mannucci Q1422 210zJHK K 3hrs arc - flat - tell
Hipp69490 Telluric 200H+K OrderSep   with Grazian Mask
Hipp69490 Telluric 210zJHK K   with Grazian Mask

Details

01:30 - Sky Flat fields in the J band - imaging
  • 2secs / 15000-10000counts 0001-0010
  • 2secs / 9000-5000counts 0011-0020
  • 2secs / <5000counts 0021-0030

02:10 - Pointing Star High Elevation (WT10_303) 0031

02:15 - Pointing Star Low Elevation (ACT0056) 0032-0034

02:21 - Point Star High Elevation (WT10_303) 0035

02:25 - Point Star near the first observing field of the night (WT10_220) - image not saved

02:28 - Persson standard BS9138 for collimation (POSITION,ACTIVE)

02:30 - Collimation - preset to Zamorani (COSMOS-cla1) acquisition script

02:37 - SCI script for Zamorani
  • Increase exposure time to 60 sec because sources too faint to see. 0036-0040
  • Cla1 field aborted because slit-mask is flipped

03:04 - preset to Zamorani (COSMOS-cla2) acquisition script 0041-0042
  • Cla2 field aborted because slit-mask is flipped

03:25 - preset to Zamorani (COSMOS-cla4) acquisition script 0043-0044
  • Cla4 field aborted because slit-mask is flipped

03:40 - preset to Mannucci (HDF) acquisition script
  • script aborted because the guide star is too far away. We did not manage to find suitable guide stars manually.

04:14 - preset to Grazian (POINT1) acquisition script 0045-0048

04:46 - starting Grazian SCI script and begin with the MOS spectroscopy 0049-0060

05:55 - repeat Grazian SCI script on the same field for the acquisition of the 2nd hour 0061-0072

07:01 - Calibration Unit on - Flat acquisition script for Grazian
  • we changed lamp from halo1 to halo2
  • halo2 OFF 0073-0075
  • halo2 ON 0076-0078

07:07 - Arc acquisition script
  • Dark Ar 0079
  • ARC Ar 0080
  • Dark Ne 0081
  • ARC Ne 0082

07:15 - Telluric star (Hipp69490) from Manucci-HDF script
  • guide star too far - abort

07:25 - Telluric star (Hipp69490) from Manucci-Q1422 script with the Grazian Slit Mask in
  • guide star not found

07:32 - Correct pointing with pointing star WT10_297 0083

07:35 - Telluric star from manucci-Q1422 (Hipp69490) script with the Grazian Slit Mask in
  • since no guide star available we pass to track and we calculate slit offsets manually because the shift-image function did not work
  • tests to put telluric in the slit 0084-0087
  • telluric in far right slit with Grazian filter setup (clear-Ordersep) 0088-0089
  • telluric in the same slit with Mannucci filter setup (clear-K) 0090-0091
  • put telluric in far left slit (Br-gam filter) 0092
  • telluric in the same slit with Mannucci filter setup 0093
  • telluric in far left slit with Grazian filter setup 0094

08:18 - Preset to Mannucci Q1422 ACQ_script
  • the guide star was not reachable and we changed, putting the one used for pre-imaging in February 0095-0098
  • MOS Aquisition application to place the sources in the slits crashes. We have to repeat.

08:55 - Starting Mannucci SCI_script for the Q1422 filed 0099-0110
  • K band - spectroscopy - script K_1
  • N1.8 / 210zJHK / clear+K/ 300secx1x12offset

10:13 - Starting Mannucci SCI_script for the Q1422 filed 0113-0124
  • K band - spectroscopy - script K_2
  • Before starting we took a through slit image to see if sources still centered 0111-0112

11:20 - Starting Mannucci SCI_script for the Q1422 filed 0125-0136
  • K band - spectroscopy - script K_3
  • skycat got stuck saving 2nd integration and before starting with the 3rd.
  • error connection to GEIRS server timed out while waiting for answer
  • it resumed on its own after a while... should be OK

12:40 - Calibration Unit on - Arc acquisition script for Mannucci Q1422
  • Dark Ar 0137
  • ARC Ar 0138
  • Dark Xe 0139
  • ARC Xe 0140

12:50 - Flat field and Darks aquisition through script DoFlats.txt
  • Series of flats with different exposure times for both Mannucci 0141-0180
  • and Grazian Masks 0181-0220

13:30 - DoDarks.txt script - End of night
  • The Darks script was interrupted - only 5sec/30sec and 2 images with 300sec darks.

-- VincenzoTest - 09 Apr 2010
Topic revision: r11 - 14 Apr 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback